OpenStudy (anonymous):

Can you check to see if I am right? The original problem is to make an equation for the line on the graph. I have provided the work and the graph, I just want to see if I am right on how to do it because I am unsure. Thanks.

OpenStudy (anonymous):

This is the graph.

OpenStudy (anonymous):

 That would make the problem, y= -1/6x+b. If I took one of the points, it would be 2=-1/6(0)+b. Doing this I would get that b = 2. In other words, y=-1/6x+2 am I right?